what is 18 divided by 1 over 6
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Same thing as



Just invert and multiply.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

what is 18 divided by 1 over 6
------
Ans: 18/(1/6) = 18*(6/1) = 108
Cheers,
Stan H.
=============